Technical Staff Blog

Last update on .

It was an interesting start to a Wednesday morning...

As you're probably already aware, there was a short blip in connectivity/functionality this morning that blocked our SSH gateway, logins and general use of the department's systems.  Our GPFS file system was overwhelmed, which resulted in a kind of traffic jam of requests ...

Last update on .

ibm-gpfs.jpg Starting at 7am on Wednesday, Nov. 9th, we will be performing rolling upgrades of our departmental GPFS NFS cluster. Each server upgrade can potentially cause a brief 90 second NFS blip for departmental linux machines. We expect the maintenance to be completed by 8am. The compute cluster linux machines and external facing services are not ...

Last update on .

network-2400px.png CIS replaced the 5th floor UPS in the switch closet this morning. Unfortunately, two of the switches failed to boot up again. CIS is working on replacing the failed switches, we will update you with an ETA once CIS provides us an update.

UPDATES:

08:02: CIS estimates it will be another hour before the ...

Last update on .

ibm-gpfs.jpg We will be upgrading GPFS from 3.5 to 4.1 on the CIFS cluster on Wednesday, Oct. 26th, between 7am and 8am.  The cifs.cs.brown.edu server will unstable during this maintenance window .  We recommend users logging out of their managed Windows machine before the 7am, Wednesday, Oct. 26th maintenance window.  

More info ...

Last update on .

All linux desktop machines will be rebooted this Tuesday, October 25th, at 5:00am. 

Please save your work before that time. If you wish to reboot before Tuesday, you may run the command:

   /local/bin/kstatus

to see if your machine is ready to reboot. If it is, ask a tstaffer to do it (by ...

Last update on .

During a routine update of our firewall there was an unexpected outage, which we are working to identify the cause of.  Most systems and connectivity have already been restored and the remaining systems will be brought back online in just a few minutes.

If you continue to have trouble after 4pm, please email problem@cs ...

Last update on .

The CIS Networking Group will be swapping out a UPS in the 5th floor networking closet on Tuesday, Oct. 25th, 6am-7am. The networking switches in the closet will loose power during this maintenance.  These switches provide networking connections to southwest section of the CIT building on the 4th and 5th.

The following rooms on the ...

Last update on .

ibm-gpfs.jpg We are currently experiencing some issues with our GPFS file system, which are causing logins to the departmental systems to hang. Updates to this blog post will be added as we debug the issue.

08:34 -  A file system fsck process appears to be hung. A support call to IBM has been initiated as we ...

Last update on .

Brown Secure Wifi Cert Update

CIS has pushed an updated certificate for the "Brown" secure wireless network.  If you're having connectivity issues with one of Brown's wifi networks, please try connecting to the "Brown-Guest" network to visit https://wifi.brown.edu/ and follow the on-screen instructions to re-estabilish your connection to the secure network(s).

For further assistance ...

Last update on .

We are about to perform an emergency failover of the department FastX gateway, which should be complete by 12:00pm. This will mean that anyone who is presently connected to FastX will lose their session(s), along with any unsaved work.

Several users reported over the weekend that the FastX gateway was refusing new connections ...

Yearly archives